U3 needs $192 million to stay in business. If it issues new common stock to raise the funds, the flotation costs will be 8%. The new issue will also require U3 to pay $280,000 in fees to its lawyers, printing cost, and other costs associated with the issue. U3 can issue stock at $25 per share. How many shares of common stock must U3 issue so that $192 million after flotation costs? Show how much of the total dollar amount of the issue will be flotation costs and how much U3 will receive after the flotation costs are paid.
Answer #1

Flotation Cost = 8%

Total Amount Raised = x

So,

x = 192,000,000 + 0.08x + 280,000

x = 192,280,000/0.92

x = $209,000,000

Number of Shares to be issued = 209,000,000/25

Number of Shares to be issued = 8,360,000